OpenStudy (anonymous):

calculate (-1+i)/(1+i) in two different ways. first, make the denominator into a real number. second, convert both numbers to polar form and use the formula for division of numbers in polar form.

OpenStudy (anonymous):

So, do you know about complex conjugates?

OpenStudy (anonymous):

If you don't have any notes about converting to polar coordinates, this is a good start: http://www.chemistrylearning.com/geometrical-to-polar-conversion/ it's pretty easy

OpenStudy (anonymous):

since it's so easy why don't I set up the first part:\[\frac{-1+i}{\ \ 1+i}\ \ \frac{1-i}{\ 1-i} = \]

OpenStudy (anonymous):

i think i know about the basics but i really can't understand what the question wants...

OpenStudy (anonymous):

I'll walk you through it. Ready?

OpenStudy (anonymous):

yeah...thx

OpenStudy (anonymous):

Can you calculate the numerator and denominator above?

OpenStudy (anonymous):

what do you mean by calculate?!calculate what?!

OpenStudy (anonymous):

(1 + i) (1 - i) = ?

OpenStudy (anonymous):

is it 2?

OpenStudy (anonymous):

Yes! (1)(1) - i^2 = 2 How about the numerator?

OpenStudy (anonymous):

(-1)(1) - i^2 = ?

OpenStudy (anonymous):

is it 2i?

OpenStudy (anonymous):

what is - i^2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

So -1 + 1 = 0. Now we have 0/1 = 0. Okay?

OpenStudy (anonymous):

oh ... my mistake ... right

OpenStudy (anonymous):

For a number a + bi, the complex conjugate is a - bi, and the product is a^2 + b^2. So multiplying by the complex conjugate of the denominator makes the denominator a real number, not complex.

OpenStudy (anonymous):

To convert to polar form, you need to get r and theta. Do you know how to do that?

OpenStudy (anonymous):

oh...got it...yeah i know about them...

OpenStudy (anonymous):

We have -1 + i and 1 + i. What is r? It's the same for both.

OpenStudy (anonymous):

square root of 2?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Do you know the formula for theta?

OpenStudy (anonymous):

tan^-1 (b/a)

OpenStudy (anonymous):

So what is theta for 1 + i?

OpenStudy (anonymous):

pi/4

OpenStudy (anonymous):

and for -1+i is 3pi/4?

OpenStudy (anonymous):

Yes. Last step.

OpenStudy (anonymous):

To divide two complex numbers we divide r for each and subtract theta

OpenStudy (anonymous):

oh thanksss...got it all...thx...:)

OpenStudy (anonymous):

I get r = 1, theta = pi, @TuringTest have I done something wrong? Please help.

OpenStudy (anonymous):

my theta is pi/2...3pi/4 - pi/4 = pi/2

OpenStudy (anonymous):

theta 1 is pi/4 theta 2 is -3pi/4 I get pi

OpenStudy (anonymous):

So I get x =1, y = 0 But we calculated the quotient is equal to 0 above.

OpenStudy (anonymous):

Oh I switched thetas.

OpenStudy (anonymous):

What do you do when converting back to x,y coord and theta is pi/2? tan is undefined.

OpenStudy (anonymous):

@eliassaab Help?

OpenStudy (anonymous):

Well, come back in a while and maybe they will have seen the pings. I'll see if I can figure it out too.

OpenStudy (anonymous):

sure ... thx ... :)

OpenStudy (anonymous):

Ah. z=r cos(theta) +i r sin(theta)

OpenStudy (anonymous):

We have 0 for the first term.

OpenStudy (anonymous):

And i for the second.

OpenStudy (anonymous):

But it still doesn't match.

OpenStudy (anonymous):

I know what I did. In multiplying the numerator (-1 + i)(1 - i) I did it wrong and forgot an imaginary part.

OpenStudy (anonymous):

am going through it again...mayb i've calculated something wrong!!!

OpenStudy (anonymous):

We should get 2i/2 = 1.

OpenStudy (anonymous):

Whew!

OpenStudy (anonymous):

I mean, equal i.

OpenStudy (anonymous):

Sound right?

OpenStudy (anonymous):

yeah...totally...got it...

OpenStudy (anonymous):

Thanks for hanging in there.

OpenStudy (anonymous):

thanks...u helped me alot...:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[ \frac{-1+i}{1+i}=\frac{(-1+i) (-1+i)}{(1+i) (-1+i)}=\frac{i^2+1-2 i}{-1+i^2}=\frac{-2 i}{-2}=i\\ \frac{-1+i}{1+i}=\frac{\sqrt{2}\, e^{\frac{i 3 \pi }{4}}}{\sqrt{2}\, e^{\frac{i \pi }{4}}}=e^{\frac{i \pi }{2}}=i \]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw